OpenStudy (anonymous):

Is the value of (1-exp(-x))/(1+exp(-x)) between (-1,1)? How?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

are you asking if that is the range?

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

\[f(x)=\frac{1-\frac{1}{e^x}}{1+\frac{1}{e^x}}\] easier to see if you multiply by \(e^x\)

OpenStudy (misty1212):

you get \[f(x)=\frac{e^x-1}{e^x+1}\]

OpenStudy (misty1212):

now it is pretty clearly always less than 1, since \(e^x>0\) for all \(x\) and the numerator is one less than the denominator

OpenStudy (anonymous):

But the min value is not -1, right?

OpenStudy (misty1212):

it cannot be \(-1\) but that is the lower limit, which is why the range is the open interval \((-1,1)\) not the closed one \([-1,1]\)

OpenStudy (anonymous):

If xtends to -inf, the value is -1 but if x tends to inf it is not 1

OpenStudy (misty1212):

why not?

OpenStudy (anonymous):

The upper limit is not one, right?

OpenStudy (misty1212):

sure it is

OpenStudy (anonymous):

But e to the power inf is infinity, so if x tends to inf, we have inf divided by inf which is not defined

OpenStudy (misty1212):

?

OpenStudy (misty1212):

you are talking about a limit right?

OpenStudy (anonymous):

Ya

OpenStudy (misty1212):

then lets go slow, because the statement "if x tends to inf, we have inf divided by inf which is not defined" makes no sense in terms of limits

OpenStudy (anonymous):

I m nr sure how the upper limit of the range is 1

OpenStudy (misty1212):

what do you make of \[\lim_{x\to \infty}\frac{x^2+1}{x^2-1}\]?

OpenStudy (anonymous):

The numerator goes to inf faster so the limit is inf

OpenStudy (anonymous):

Or Do we divide by x n see

OpenStudy (misty1212):

oh no i see the confusion \[x^2+1\] goes to infinity just as fast as \(x^2-1\) the number out at the end makes no difference at all

OpenStudy (anonymous):

So it tends to 1

OpenStudy (misty1212):

you remember in some class, maybe called "pre calculus" finding the horizontal asymptote of a rational function?

OpenStudy (anonymous):

Ya

OpenStudy (misty1212):

that is what you are doing exactly if the degrees are the same, it is the ratio of the leading coefficients you do not say "it is infinity over infinity therefore undefined"

OpenStudy (anonymous):

I m still not getting why the upper limit of my expression shd be 1, cn u pls explain that

OpenStudy (misty1212):

of the expression \[\frac{e^x-1}{e^x-1}\]?

OpenStudy (anonymous):

Ya

OpenStudy (misty1212):

ooops i made a typo there, should be \[\frac{e^x-1}{e^x+1}\]

OpenStudy (misty1212):

ok we agree that the numerator is one less then the denominator right? and therefore it can never be one, because a fraction is only one if the numerator and denominator are the same

OpenStudy (anonymous):

Ya

OpenStudy (misty1212):

and also since the numerator is less than the denominator, this is always smaller than 1, not larger right?

OpenStudy (anonymous):

Right

OpenStudy (misty1212):

so let me ask you a question it can never be one it is always smaller than one what do you think the limit might be? \(\frac{1}{2}\)?

OpenStudy (misty1212):

\(e^x\) grows very very quickly, much faster than \(x^2\) or in fact any power of \(x\)

OpenStudy (anonymous):

Ok

OpenStudy (misty1212):

already \(e^{10}\) is \(22026\) rounded so you would have \[\frac{22025}{22027}\] which is pretty close to 1, with \(x=10\) only

OpenStudy (anonymous):

I see

OpenStudy (misty1212):

imagine what it would look like if \(x=100\)

OpenStudy (anonymous):

Can u tell me what kind of function will have a range between 1 and infinity

OpenStudy (misty1212):

i can make up one sure

OpenStudy (misty1212):

how about \[f(x)=\sqrt{x}+1\]

OpenStudy (misty1212):

or \[f(x)=x^2+1\]

OpenStudy (anonymous):

I see that

OpenStudy (misty1212):

those would have range \([1,\infty)\) if you want \((1,\infty)\) open you could use for example \[f(x)=e^x+1\]

OpenStudy (anonymous):

I was looking for functions close to logit

OpenStudy (misty1212):

ok wise guy

OpenStudy (perl):

\(\color{blue}{\text{Originally Posted by}}\) @misty1212 now it is pretty clearly always less than 1, since \(e^x>0\) for all \(x\) and the numerator is one less than the denominator \(\color{blue}{\text{End of Quote}}\) The numerator* is 2 less than the denominator

OpenStudy (misty1212):

you must be bored

OpenStudy (anonymous):

No way

OpenStudy (anonymous):

U explain so well

OpenStudy (anonymous):

R u prof of math?